What is the amount of power applied by the wind glass to a 20kg anchor if
it is lifted 2-m high within 5s?
1
Expert's answer
2020-10-30T13:19:49-0400

By definition, the power is:


"P = Fv"

where "v = \dfrac{2m}{5s} = \dfrac{2}{5}m/s" is the speed of anchor and "F = mg = 20kg\cdot 9.81m/s^2 = 196.2N" is the force applied to it. Thus, obtain:


"P = Fv = \dfrac{2}{5}\cdot 196.2 = 78.48W"

Answer. 78.48 W.